### v1

# FMPlug: Plug-In Foundation Flow-Matching Priors for Inverse Problems

Abstract
We present FMPlug, a novel plug-in framework that enhances foundation flow-matching (FM) priors for solving ill-posed inverse problems. Unlike traditional approaches that rely on domain-specific or untrained priors, FMPlug smartly leverages two simple but powerful insights: the similarity between observed and desired objects and the Gaussianity of generative flows. By introducing a time-adaptive warm-up strategy and sharp Gaussianity regularization, FMPlug unlocks the true potential of domain-agnostic foundation models. Our method beats state-of-the-art methods that use foundation FM priors by significant margins, on image super-resolution and Gaussian deblurring.
Index Terms—  Flow matching, generative models, inverse problems, image priors, image restoration
1 Introduction
Inverse problems (IPs) are prevalent in many fields, such as medical imaging, remote sensing, and computer vision. IPs aim to recover an unknown object $\bm{x}$ from its measurement $\bm{y}\approx\mathcal{A}(\bm{x})$, where the forward operator $\mathcal{A}(\cdot)$ models the measurement process and the approximation sign $\approx$ accounts for potential modeling errors and measurement noise. Due to insufficient measurement and/or the approximate relationship in $\bm{y}\approx\mathcal{A}(\bm{x})$, in practice $\bm{x}$ is typically not uniquely recoverable from $\bm{y}$ alone, i.e., ill-posedness. So, to obtain a reliable and meaningful solution for IPs, it is important to incorporate prior knowledge of $\bm{x}$.
Traditional ideas for solving IPs rely on optimization formulations and associated numerical optimization methods, often motivated under the Maximum A Posteriori (MAP) estimation principle:
$\min\nolimits_{\bm{x}}\;\ell(\bm{y},\mathcal{A}(\bm{x}))+\Omega(\bm{x}).$ (1.1)
Here, minimizing data fitting loss $\ell(\bm{y},\mathcal{A}(\bm{x}))$ encourages $\bm{y}\approx\mathcal{A}(\bm{x})$, and the regularization term $\Omega(\bm{x})$ encodes prior knowledge of ideal solutions to resolve ambiguities and hence mitigate potential ill-posedness. Advances in deep learning (DL) have revolutionized IP solving. Different DL-based approaches for IPs operate with different data-knowledge tradeoffs. For example, supervised approaches take paired datasets $\left\{\bm{y}_{i},\bm{x}_{i}\right\}_{i=1}^{N}$ and directly learn the inverse mapping, with or without using $\mathcal{A}$ [1, 2, 3]; alternative, data-driven priors learned from object-only datasets $\left\{\bm{x}_{i}\right\}_{i=1}^{N}$ can be integrated with Eq. 1.1 [4, 5, 6]; strikingly, untrained DL models themselves can serve as plug-in priors for Eq. 1.1, without any extra data [7, 8, 9, 10, 11, 12]. [1] give comprehensive reviews of these ideas.
In this paper, we focus on solving IPs with pretrained deep generative priors (DGPs) [4]. Compared to supervised approaches that perform task-specific training, these approaches enjoy great flexibility, as DGPs can be plugged in and reused for different tasks related to the same family of objects. Among the different DGPs, we are most interested in those based on diffusion models (DM), and the more general flow-matching (FM) models [13], as they are the backbone of state-of-the-art (SOTA) generative models in several domains [14, 15, 16].
Numerous recent works target solving IPs with pretrained DM/FM priors [5]. Although promising, most of them are based on domain-specific DM/FM priors, e.g., trained on the FFHQ dataset for human faces and the LSUN bedrooms dataset for bedroom scenes. This limits the practicality of these methods, as domain-specific DM/FM models are not always readily available. On the other hand, the emergence of domain-agnostic foundation DM/FM models, such as Stable Diffusion [15] and Flux [14] for images, seems to obsolete domain-specific developments; [17, 18, 19] propose such ideas. However, the reported performance clearly lags behind those obtained with domain-specific FM/DM priors, and even behind those obtained with untrained priors; see Section 2.3. This should not be surprising, as in terms of solving any particular IP, foundation priors are considerably weaker than domain-specific ones.
In this paper, we describe our ongoing efforts to close the performance gap. We focus on image restoration and take advantage of two things to strengthen the image prior: (1) the similarity of $\bm{x}$ and $\bm{y}$, which motivates us to develop a learnable warm-up strategy; and (2) Gaussianity in DM/FM priors, suggesting Gaussianity regularization. By incorporating these two simple changes, we demonstrate convincing performance gains over prior work based on foundation FM priors.
2 Technical background & related Work
2.1 Flow matching (FM)
Flow Matching (FM) models are an emerging class of deep generative models [13]. They learn a continuous flow to transform a prior distribution $p_{0}(\bm{z})$ into a target distribution $p_{1}(\bm{x})$—in the same spirit of continuous normalizing flow (CNF), described by an ordinary differential equation (ODE)
$d\bm{z}=\bm{v}(\bm{z},t)\ dt.$ (2.1)
Whereas CNF focuses on the density path induced by the flow and performs the maximum likelihood estimation, FM tries to learn a parametrized velocity field $\bm{v}_{\bm{\theta}}(\bm{z},t)$ to match the one associated with the desired flow. To generate new samples after training, one simply samples $\bm{z}_{0}\sim p_{0}(\bm{z})$ and numerically solves the learned ODE induced by $\bm{v}_{\bm{\theta}}(\bm{z},t)$ from $t=0$ to $t=1$, to produce a sample $\bm{z}_{1}\sim p_{1}(\bm{x})$.
For tractability, in practice, FM matches the conditional velocity field: for each training point $\bm{x}$, a simple conditional probability path $p_{t}(\bm{z}|\bm{x})$, e.g., induced by a linear interpolation $\bm{z}_{t}=t\bm{x}+(1-t)\bm{z}_{0}$, is defined. The model is then trained to learn the known vector field of this conditional path
$\min\nolimits_{\bm{\theta}}\;\mathbb{E}_{\bm{x},\bm{z}_{0},t}\left\|\bm{v}_{\theta}(\bm{z}_{t},t)-\bm{u}(\bm{z}_{t},t|\bm{x})\right\|^{2}.$ (2.2)
DMs based on probability flow ODEs can also be interpreted as FMs, although they match the score functions induced by the chosen probability path, not the vector field [13]. So, FM can be viewed as a general generative framework that covers DM as well besides other algorithmic possibilities.
2.2 Pretrained FM priors for IPs
Recent methods that use pretrained FM priors for solving IPs can be classified into two families: (1) The interleaving approach interleaves the generation steps (i.e., numerical integration steps) with gradient steps toward feasibility (i.e., making $\bm{x}$ satisfies that $\bm{y}\approx\mathcal{A}(\bm{x})$[20, 17, 18]; see a sample algorithm in Algorithm 1. Despite the simplicity and empirical effectiveness on simple IPs [20, 17, 18], these methods might not converge or return an $\bm{x}$ that respects the data prior (i.e., manifold feasibility) and/or satisfies the measurement constraint $\bm{y}\approx\mathcal{A}(\bm{x})$ (i.e., measurement feasibility); and (2) the plug-in approach views the entire generation process (lines 4–8 in Algorithm 2) as a function $\mathcal{G}_{\bm{\theta}}$ that maps any source sample to a target sample, and plugs the prior into Eq. 1.1 to obtain a unified optimization formulation [19, 6, 21] ($\circ$: function composition):
$\bm{z}^{\ast}\in\operatorname*{arg\,min}\nolimits_{\bm{z}}\;\mathcal{L}(\bm{z})\doteq\ell\pqty{\bm{y},\mathcal{A}\circ\mathcal{G}_{\bm{\theta}}(\bm{z})}+\Omega\circ\mathcal{G}_{\bm{\theta}}(\bm{z});$ (2.3)
see Algorithm 2 for a sample algorithm. The estimated object is then $\mathcal{G}_{\bm{\theta}}(\bm{z}^{\ast})$. Here, the generator $\mathcal{G}_{\bm{\theta}}$ is fixed and the output $\mathcal{G}_{\bm{\theta}}(\bm{z})$ naturally satisfies the manifold feasibility. In addition, global optimization of $\mathcal{L}(\bm{z})$ forces small ($)\ell\pqty{\bm{y},\mathcal{A}\circ\mathcal{G}_{\bm{\theta}}(\bm{z})}$, and hence $\bm{y}\approx\mathcal{A}\circ\mathcal{G}_{\bm{\theta}}(\bm{z})$, i.e., measurement feasibility.
Algorithm 1 A sample algorithm of the interleaving approach
1:ODE steps $T$, measurement $\bm{y}$, forward model $\mathcal{A}$
2:Initialize $\bm{z}_{0}\sim\mathcal{N}(\bm{0},\bm{I})$
3:for $i=0$ to $T-1$ do
4:  $t_{i}\leftarrow i/T$
5:  $\bm{v}\leftarrow\bm{v}_{\bm{\theta}}(\bm{z}_{i},t_{i})$triangleright learned velocity
6:  $\bm{z}^{\prime}_{i+1}\leftarrow\bm{z}_{i}+1/T\cdot\bm{v}$triangleright discrete integration
7:  $\bm{z}_{i+1}\leftarrow$ $(\bm{y},\mathcal{A})$-driven update of $\bm{z}^{\prime}_{i+1}$triangleright reducing $\ell(\bm{y},\mathcal{A}(\bm{z}))$ starting from $\bm{z}^{\prime}_{i+1}$
8:end for
9:Estimated $\widehat{\bm{x}}$
Algorithm 2 A sample algorithm of the plug-in approach
1:Total iterations $E$, ODE steps $T$, measurement $\bm{y}$, forward model $\mathcal{A}$
2:Initialize $\bm{z}^{(0)}\sim\mathcal{N}(\bm{0},\bm{I})$
3:for $e=0$ to $E-1$ do
4:  $\bm{z}_{0}\leftarrow\bm{z}^{(e)}$
5:  for $i=0$ to $T-1$ do $\triangleright$ whole integration path
6:   $t_{i}\leftarrow i/T$
7:   $\bm{v}\leftarrow\bm{v}_{\bm{\theta}}(\bm{z}_{i},t_{i})$
8:   $\bm{z}_{i+1}\leftarrow\bm{z}_{i}+1/T\cdot\bm{v}$
9:  end for
10:  $\bm{z}^{(e+1)}\leftarrow\bm{z}^{(e)}-\eta^{(e)}\cdot\nabla_{\bm{z}}\mathcal{L}(\bm{z}_{i+1})$triangleright reducing loss
11:end for
12:Estimation $\widehat{\bm{x}}=\mathcal{G}_{\bm{\theta}}(\bm{z}^{(E-1)})$
2.3 Foundation FM priors for IPs
Table 1: A comparison between foundation, domain-specific, and untrained priors for Gaussian deblurring on AFHQ (resolution: $\times 256$ 256; DS: domain-specific; FD: foundation). Bold: best, & underline: second best, for each metric/column
[…]
DIP 32.9470 0.9075 0.1873 0.5024
D-Flow (DS) 30.2512 0.8553 0.1183 0.6536
D-Flow (FD) 30.0975 0.8861 0.2270 0.5404
The availability of large-scale training data has enabled the development of foundation generative models for several domains, most of them based on FM (or DM) models, e.g., stable diffusion (v3) [15] and FLUX [14] for images, Sora [22] and Veo for videos, Cosmos world model [16]. So, recent IP methods based on pretrained FM priors have started to shift from domain-specific priors to these foundation priors.
However, while these foundation priors only constrain the object to be physically meaningful (e.g., the object being a natural image), domain-specific priors provide much more semantic and perhaps structural information about the object (e.g., the object being a facial or brain MRI image), i.e., foundation priors alone are considerably weaker than domain-specific priors. In fact, untrained priors such as deep image priors may be powerful enough to promote physically meaningful reconstruction [7, 8, 9, 10, 11]. A quick comparison summarized in Table 1 confirms our intuition: recent IP methods with foundation priors perform worst compared to domain-specific and even untrained priors.
To strengthen the object prior, [19] assumes similarity of $\bm{x}$ and $\bm{y}$—valid for typical image restoration problems, and initializes Eq. 2.3 with
$\bm{z}_{0}=\sqrt{\alpha}\bm{y}_{0}+\sqrt{1-\alpha}\bm{z}\quad\text{where}\;\bm{z}\sim\mathcal{N}(\bm{0},\bm{I}),$ (2.4)
[…]
3 Method
We focus on image restoration with foundation FM priors, and aim to strengthen the image prior based on (1) the similarity of $\bm{x}$ and $\bm{y}$, and (2) Gaussianity in FM priors, in the same vein as [19] but improving upon theirs. Adding text prompts as in [17] is presumably helpful, but the reliable generation of high-quality prompts for degraded images seems tricky; we do not pursue it here for leave it for future work.
3.1 Time-dependent warm-up strategy
In the standard FM setting, the source distribution $\bm{z}_{0}\sim\mathcal{N}(\bm{0},\bm{I})$, whereas the initialized $\bm{z}_{0}$ in Eq. 2.4 has a distribution $\mathcal{N}(\sqrt{\alpha}\bm{y}_{0},(1-\alpha)\bm{I})$. One might not worry about this distribution mismatch, as both are supported on the entire ambient space anyway. But finite-sample training in practice causes the gap: due to the famous concentration of measure phenomenon [23], virtually all training samples drawn from $\mathcal{N}(\bm{0},\bm{I})$ come from an ultra-thin shell around $\mathbb{S}^{d-1}(\bm{0},\sqrt{d})$, a sphere centered at $\bm{0}$ and with a radius $\sqrt{d}$: $\mathbb{P}[|\norm{\bm{z}}_{2}-\sqrt{d}|\geq t]\leq 2e^{-ct^{2}}$ for a universal constant $c>0$. If we write that ultra-thin shell spanned by the finite training set as $\mathcal{S}$, the generation function $\mathcal{G}_{\bm{\theta}}$ is effectively trained on inputs from the domain $\mathcal{S}$, not the entire ambient space: the behavior of $\mathcal{G}_{\bm{\theta}}$ on $\mathcal{S}^{c}$ is largely undetermined. Now, samples from $\mathcal{N}(\sqrt{\alpha}\bm{y}_{0},(1-\alpha)\bm{I})$ concentrate around another ultra-thin shell around $\mathbb{S}^{d-1}(\sqrt{\alpha}\bm{y}_{0},\sqrt{(1-\alpha)d})$, which has only a negligibly small intersection with $\mathcal{S}$. So, the initialization in Eq. 2.4 lies in $\mathcal{S}^{c}$ with a very high probability. Given that the behavior of $\mathcal{G}_{\bm{\theta}}$ can be wild on $\mathcal{S}^{c}$, this initialization strategy seems problematic.
We take a different route. We know that a typical flow of FM models takes the form:
$\bm{z}_{t}=\alpha_{t}\bm{x}+\beta_{t}\bm{z}\quad\text{where}\;\bm{z}\sim\mathcal{N}(\bm{0},\bm{I}).$ (3.1)
Now, when $\bm{x}$ and $\bm{y}$ are close, $\bm{x}=\bm{y}+\bm{\varepsilon}$ for some small $\bm{\varepsilon}$. So, we can write the flow as
$\bm{z}_{t}=\alpha_{t}(\bm{y}+\bm{\varepsilon})+\beta_{t}\bm{z}\quad\text{where}\;\bm{z}\sim\mathcal{N}(\bm{0},\bm{I})$ (3.2)
for unknown $\bm{\varepsilon}$. When $\alpha_{t}$ is sufficiently small—i.e., we are sufficiently close to $t=0$ in the flow, $\alpha_{t}\bm{\varepsilon}$ can be negligibly small, leading to the approximate flow
$\bm{z}_{t}\approx\alpha_{t}\bm{y}+\beta_{t}\bm{z}\quad\text{where}\;\bm{z}\sim\mathcal{N}(\bm{0},\bm{I}).$ (3.3)
In practice, we do not know how small $\alpha_{t}$ should be, so we leave it learnable. This leads to the following formulation
$)).\displaystyle\boxed{\min\nolimits_{\bm{z},t\in[0,1]}\;\ell\pqty{\bm{y},\mathcal{A}\circ\mathcal{G}_{\bm{\theta}}\pqty{\alpha_{t}\bm{y}+\beta_{t}\bm{z},t}}}.$ (3.4)
Note that here we overload the notation of $\mathcal{G}_{\bm{\theta}}$ as $\mathcal{G}_{\bm{\theta}}:\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}$—the second input is the current $t$ on the path (the notation we use in Eq. 2.3 assumes $t=0$). In other words, due to the closeness of $\bm{x}$ and $\bm{y}$, we do not need to start from scratch, i.e., a random sample from the source distribution; instead, we plug $\bm{y}$ into an appropriate—to be learned—time point of the flow to create a shortcut.
Our formulation in Eq. 3.4 can be easily generalized to latent FM models. Moreover, it is not only grounded in theory and effective in practice (see Section 4), but also speeds up learning as $t>0$ implies shorter flows—although improving the speed is not our current focus.
Due to the approximation errors in matching the ideal flow during FM training, as well as when approximating Eq. 3.2 using Eq. 3.3, the distribution of $\bm{z}_{t}$ could be slightly off the ideal distribution. To rectify this, we perform a lightweight variance calibration in our implementation:
$\hat{\bm{z}}_{t}=\sqrt{\mathrm{Var}(Z_{t})/\mathrm{Var}(\bm{z}_{t})}\cdot\bm{z}_{t},$ (3.5)
where $\mathrm{Var}(Z_{t})$ is the scalar-variable variance across all dimensions and the generation paths of 512512512 samples, and $\mathrm{Var}(\bm{z}_{t})$ scalar-variable variance for $\bm{z}_{t}$ across all dimensions.
3.2 Gaussianity regularization
Fig. 1: Plot of the function $h(\bm{z}_{0})$ (after a change of variable $u=\norm{\bm{z}_{0}}_{2}^{2}$). An ideal regularization function should blow up sharply away from the narrow concentration region in orange to promote Gaussianity effectively.
If $\bm{z}_{0}\sim\mathcal{N}(\bm{0},\bm{I})$, $\norm{\bm{z}_{0}}_{2}^{2}\sim\chi^{2}(d)$ and the negative log-likelihood is $h(\bm{z}_{0})=-(d/2-1)\log\norm{\bm{z}_{0}}_{2}^{2}+\norm{\bm{z}_{0}}_{2}^{2}/2+C$ for some constant $C$ independent of $\bm{z}_{0}$. [19] promotes the Gaussianity of $\bm{z}_{0}$ by regularizing $h(\bm{z}_{0})$. While $h(\bm{z}_{0})$ is minimized at any $\bm{z}_{0}$ satisfies $\norm{\bm{z}_{0}}_{2}^{2}=d-2$, away from this value the function changes slowly; see Fig. 1. This is problematic, as $\norm{\bm{z}_{0}}_{2}^{2}$ should concentrate sharply around $d$ due to concentration of measure and thus only functions that blow up quickly away from the $\norm{\bm{z}_{0}}_{2}^{2}=d$ level can effectively promote the Gaussianity of $\bm{z}_{0}$.
For our formulation in Eq. 3.4, we hope to promote the Gaussianity of $\bm{z}$. To enforce the sharp concentration of $\bm{z}$, we project $\bm{z}$ onto the sphere $\mathbb{S}^{d-1}(\bm{0},\sqrt{d})$ in each step by
$\bm{z}^{\prime}=\sqrt{d}\bm{z}/\|\bm{z}\|_{2}.$ (3.6)
We are not the first to use the more precise spherical constraint or regularization in promoting Gaussianity in the DM/FM literature; see, e.g., [24].
4 Experiment
Table 2: x4 Super Resolution with additive Gaussian noise ($\sigma=0.03$). (Bold: best, underline: second best)
DIV2K [25] AFHQ [26]
[…]
DIP 24.896 0.613 0.482 0.305 28.275 0.697 0.466 0.284
FlowChef 24.439 0.677 0.444 0.630 28.444 0.779 0.404 0.568
FlowDPS 24.284 0.657 0.461 0.432 28.507 0.743 0.445 0.293
D-Flow 24.467 0.684 0.488 0.387 27.596 0.709 0.518 0.296
FMPlug-W 26.047 0.762 0.369 0.410 30.234 0.813 0.358 0.367
FMPlug-W-R 26.144 0.769 0.355 0.497 30.449 0.821 0.340 0.473
Table 3: Gaussian Deblur with additive Gaussian noise ($\sigma=0.03$). (Bold: best, underline: second best)
DIV2K [25] AFHQ [26]
[…]
DIP 24.659 0.658 0.469 0.367 28.169 0.680 0.472 0.283
FlowChef 20.248 0.474 0.629 0.236 24.366 0.661 0.533 0.298
FlowDPS 19.900 0.436 0.615 0.190 24.612 0.626 0.532 0.155
D-Flow 24.798 0.683 0.390 0.464 28.955 0.752 0.458 0.316
FMPlug-W 25.995 0.749 0.387 0.413 30.315 0.805 0.378 0.246
FMPlug-W-R 26.125 0.757 0.378 0.412 30.322 0.808 0.372 0.248
We test our FMPlug method against other SOTA methods on two linear IPs: super-resolution and Gaussian debluring.
Datasets, tasks, and evaluation metrics  Following [17], we use two datasets: 252525 DIV2K  [25] validation images and 252525 AFHQ cat  [26] images. We set the image resolution to $\times 512$ 512 by resizing and cropping the original. We consider two tasks: i) 4$\times$ super-resolution (SR) from $\times 128$ 128 to $\times 512$ 512; ii) Gaussian deblurring (GD) with a kernel size of 616161 and standard deviation of 3.03.03.0. Both tasks are added with Gaussian noise $\sigma=0.03$. For metrics, we use PSNR and SSIM for pixel-level difference, and LPIPS for perceptual difference, and CLIPIQA, a no-reference quality metric.
[…]
No text-prompt For a fair comparison, we disable text prompt and classifier-free guidance for all methods. It is also not clear how reliably we can generate pertinent text prompts for degraded images.
ObservationsTable 2 and Table 3 summarize the quantitative results. We can observe that: (1) Our FMPlug-W-R is the overall winner by all metrics but CLIPIQA, a no-reference metric. While on the relatively easy SR task the performance gaps between different methods seem minor, on GD the gaps are significant: FlowChef and FlowDPS without prompt guidance lag behind even the untrained DIP by large margins and generate visually blurry and oversmooth images as shown in Figs. 2, 5, 4 and 3, suggesting the general struggle of interleaving methods to ensure simultaneous measurement and manifold feasibility; (2) For plug-in methods, our FMPlug-W improves upon D-Flow, by considerable margins based on all metrics but CLIPIQA, showing the solid advantage of our warm-up strategy in Eq. 3.4 over their initialization strategy in Eq. 2.4; and (3) FMPlug-W-R further improves PSNR and SSIM slightly over FMPlug-W, with the largest improvement seen in CLIPIQA, showing stronger visual quality. This confirms the benefits brought about by the precise Gaussianity regularization in Eq. 3.6.
(a) Measurement
(b) Ground Truth
(c) DIP
(d) FlowChef
(e) FlowDPS
(f) D-Flow
(g) FMPlug-W
(h) FMPlug-W-R
Fig. 2: Qualitative results for 4×4$\times$4 × super-resolution on the AFHQ dataset.
(a) Measurement
(b) Ground Truth
(c) DIP
(d) FlowChef
(e) FlowDPS
(f) D-Flow
(g) FMPlug-W
(h) FMPlug-W-R
Fig. 3: Qualitative results for 4×4$\times$4 × super-resolution on the DIV2K dataset.
(a) Measurement
(b) Ground Truth
(c) DIP
(d) FlowChef
(e) FlowDPS
(f) D-Flow
(g) FMPlug-W
(h) FMPlug-W-R
Fig. 4: Qualitative results for Gaussian deblurring on the AFHQ dataset.
(a) Measurement
(b) Ground Truth
(c) DIP
(d) FlowChef
(e) FlowDPS
(f) D-Flow
(g) FMPlug-W
(h) FMPlug-W-R
Fig. 5: Qualitative results for Gaussian deblurring on DIV2K.

### v2

@@
 So, to obtain a reliable and meaningful solution for IPs, it is important to incorporate prior knowledge of $\bm{x}$.
-Traditional ideas for solving IPs rely on optimization formulations and associated numerical optimization methods, often motivated under the Maximum A Posteriori (MAP) estimation principle:
+Traditional ideas for solving IPs rely on optimization formulations, often motivated under the Maximum A Posteriori (MAP) estimation principle:
 $\min\nolimits_{\bm{x}}\;\ell(\bm{y},\mathcal{A}(\bm{x}))+\Omega(\bm{x}).$ (1.1)
 Here, minimizing data fitting loss $\ell(\bm{y},\mathcal{A}(\bm{x}))$ encourages $\bm{y}\approx\mathcal{A}(\bm{x})$, and the regularization term $\Omega(\bm{x})$ encodes prior knowledge of ideal solutions to resolve ambiguities and hence mitigate potential ill-posedness.
+The optimization problems are often solved by gradient-based iterative methods.
 Advances in deep learning (DL) have revolutionized IP solving.
-Different DL-based approaches for IPs operate with different data-knowledge tradeoffs.
-For example, supervised approaches take paired datasets $\left\{\bm{y}_{i},\bm{x}_{i}\right\}_{i=1}^{N}$ and directly learn the inverse mapping, with or without using $\mathcal{A}$ [1, 2, 3]; alternative, data-driven priors learned from object-only datasets $\left\{\bm{x}_{i}\right\}_{i=1}^{N}$ can be integrated with Eq.
-1.1 [4, 5, 6]; strikingly, untrained DL models themselves can serve as plug-in priors for Eq.
-1.1, without any extra data [7, 8, 9, 10, 11, 12].
-[1] give comprehensive reviews of these ideas.
-In this paper, we focus on solving IPs with pretrained deep generative priors (DGPs) [4].
+Different DL-based approaches to IPs operate with different data-knowledge tradeoffs.
+For example, supervised approaches take paired datasets $\left\{\bm{y}_{i},\bm{x}_{i}\right\}_{i=1}^{N}$ and directly learn the inverse mapping $\bm{y}\mapsto\bm{x}$, with or without using $\mathcal{A}$ [Ongie2020DeepLearning, monga2021algorithm, zhang2024wrong]; alternatively, data-driven priors learned from object-only datasets $\left\{\bm{x}_{i}\right\}_{i=1}^{N}$ can be integrated with Eq.
+1.1 [oliviero2025generative, daras2024surveydiffusionmodelsinverse, Wang2024DMPlugAP]; strikingly, untrained DL models themselves can serve as plug-in priors for Eq.
+1.1, without any extra data [alkhouri2025understanding, wang2023early, li_deep_2023, zhuang_blind_2023, zhuang_practical_2023, li2021self].
+[Ongie2020DeepLearning] give comprehensive reviews of these ideas.
+In this paper, we focus on solving IPs with pretrained deep generative priors (DGPs) [oliviero2025generative].
 Compared to supervised approaches that perform task-specific training, these approaches enjoy great flexibility, as DGPs can be plugged in and reused for different tasks related to the same family of objects.
-Among the different DGPs, we are most interested in those based on diffusion models (DM), and the more general flow-matching (FM) models [13], as they are the backbone of state-of-the-art (SOTA) generative models in several domains [14, 15, 16].
-Numerous recent works target solving IPs with pretrained DM/FM priors [5].
+Among the different DGPs, we are most interested in those based on diffusion models (DMs), and the more general flow-matching (FM) models [lipman2024flow], as they are the backbone of state-of-the-art (SOTA) generative models in several domains [labs2025flux1kontextflowmatching, esser2024scalingrectifiedflowtransformers, Agarwal2025Cosmos].
+Numerous recent works target solving IPs with pretrained DM/FM priors [daras2024surveydiffusionmodelsinverse].
 Although promising, most of them are based on domain-specific DM/FM priors, e.g., trained on the FFHQ dataset for human faces and the LSUN bedrooms dataset for bedroom scenes.
 This limits the practicality of these methods, as domain-specific DM/FM models are not always readily available.
-On the other hand, the emergence of domain-agnostic foundation DM/FM models, such as Stable Diffusion [15] and Flux [14] for images, seems to obsolete domain-specific developments; [17, 18, 19] propose such ideas.
+On the other hand, the emergence of domain-agnostic foundation DM/FM models, such as Stable Diffusion [esser2024scalingrectifiedflowtransformers] and Flux [labs2025flux1kontextflowmatching] for images, seems to obsolete domain-specific developments; [kim2025flowdpsflowdrivenposteriorsampling, patel2024steering, ben2024d] propose such ideas.
 However, the reported performance clearly lags behind those obtained with domain-specific FM/DM priors, and even behind those obtained with untrained priors; see Section 2.3.
-This should not be surprising, as in terms of solving any particular IP, foundation priors are considerably weaker than domain-specific ones.
+This should not be surprising, as in terms of solving any particular IP, foundation priors are considerably weaker than domain-specific ones and not clearly stronger than untrained priors.
 In this paper, we describe our ongoing efforts to close the performance gap.
@@
 2.1 Flow matching (FM)
-Flow Matching (FM) models are an emerging class of deep generative models [13].
+Flow Matching (FM) models are an emerging class of deep generative models [lipman2024flow].
 They learn a continuous flow to transform a prior distribution $p_{0}(\bm{z})$ into a target distribution $p_{1}(\bm{x})$—in the same spirit of continuous normalizing flow (CNF), described by an ordinary differential equation (ODE)
@@
 $\min\nolimits_{\bm{\theta}}\;\mathbb{E}_{\bm{x},\bm{z}_{0},t}\left\|\bm{v}_{\theta}(\bm{z}_{t},t)-\bm{u}(\bm{z}_{t},t|\bm{x})\right\|^{2}.$ (2.2)
-DMs based on probability flow ODEs can also be interpreted as FMs, although they match the score functions induced by the chosen probability path, not the vector field [13].
+DMs based on probability flow ODEs can also be interpreted as FMs, although they match the score functions induced by the chosen probability path, not the vector field [lipman2024flow].
 So, FM can be viewed as a general generative framework that covers DM as well besides other algorithmic possibilities.
 2.2 Pretrained FM priors for IPs
-Recent methods that use pretrained FM priors for solving IPs can be classified into two families: (1) The interleaving approach interleaves the generation steps (i.e., numerical integration steps) with gradient steps toward feasibility (i.e., making $\bm{x}$ satisfies that $\bm{y}\approx\mathcal{A}(\bm{x})$[20, 17, 18]; see a sample algorithm in Algorithm 1.
-Despite the simplicity and empirical effectiveness on simple IPs [20, 17, 18], these methods might not converge or return an $\bm{x}$ that respects the data prior (i.e., manifold feasibility) and/or satisfies the measurement constraint $\bm{y}\approx\mathcal{A}(\bm{x})$ (i.e., measurement feasibility); and (2) the plug-in approach views the entire generation process (lines 4–8 in Algorithm 2) as a function $\mathcal{G}_{\bm{\theta}}$ that maps any source sample to a target sample, and plugs the prior into Eq.
-1.1 to obtain a unified optimization formulation [19, 6, 21] ($\circ$: function composition):
-$\bm{z}^{\ast}\in\operatorname*{arg\,min}\nolimits_{\bm{z}}\;\mathcal{L}(\bm{z})\doteq\ell\pqty{\bm{y},\mathcal{A}\circ\mathcal{G}_{\bm{\theta}}(\bm{z})}+\Omega\circ\mathcal{G}_{\bm{\theta}}(\bm{z});$ (2.3)
-see Algorithm 2 for a sample algorithm.
-The estimated object is then $\mathcal{G}_{\bm{\theta}}(\bm{z}^{\ast})$.
-Here, the generator $\mathcal{G}_{\bm{\theta}}$ is fixed and the output $\mathcal{G}_{\bm{\theta}}(\bm{z})$ naturally satisfies the manifold feasibility.
-In addition, global optimization of $\mathcal{L}(\bm{z})$ forces small ($)\ell\pqty{\bm{y},\mathcal{A}\circ\mathcal{G}_{\bm{\theta}}(\bm{z})}$, and hence $\bm{y}\approx\mathcal{A}\circ\mathcal{G}_{\bm{\theta}}(\bm{z})$, i.e., measurement feasibility.
+Recent methods that use pretrained FM priors for solving IPs can be classified into two families: (1) The interleaving approach interleaves the generation steps (i.e., numerical integration steps) with gradient steps toward feasibility (i.e., making $\bm{x}$ satisfies that $\bm{y}\approx\mathcal{A}(\bm{x})$[pokle2023training, kim2025flowdpsflowdrivenposteriorsampling, patel2024steering, martin2025pnpflowplugandplayimagerestoration]; see a sample algorithm in Algorithm 1.
+Despite the simplicity and empirical effectiveness on simple IPs, these methods might not converge or return an $\bm{x}$ that respects the data prior (i.e., manifold feasibility) and/or satisfies the measurement constraint $\bm{y}\approx\mathcal{A}(\bm{x})$ (i.e., measurement feasibility); and (2) the plug-in approach views the entire generation process (lines 4–8 in Algorithm 2) as a function $\mathcal{G}_{\bm{\theta}}$ that maps any source sample to a target sample, and plugs the prior into Eq.
+1.1 to obtain a unified optimization formulation [ben2024d, Wang2024DMPlugAP, wang2025temporal] ($\circ$: function composition):
+[…]
+[…]
 Algorithm 1 A sample algorithm of the interleaving approach
@@
 2:Initialize $\bm{z}_{0}\sim\mathcal{N}(\bm{0},\bm{I})$
-3:for $i=0$ to $T-1$ do
+3:for $i=0$ to T1T-1 do
 4:  $t_{i}\leftarrow i/T$
-5:  $\bm{v}\leftarrow\bm{v}_{\bm{\theta}}(\bm{z}_{i},t_{i})$triangleright learned velocity
-6:  $\bm{z}^{\prime}_{i+1}\leftarrow\bm{z}_{i}+1/T\cdot\bm{v}$triangleright discrete integration
-7:  $\bm{z}_{i+1}\leftarrow$ $(\bm{y},\mathcal{A})$-driven update of $\bm{z}^{\prime}_{i+1}$triangleright reducing $\ell(\bm{y},\mathcal{A}(\bm{z}))$ starting from $\bm{z}^{\prime}_{i+1}$
+5:  $\bm{v}\leftarrow\bm{v}_{\bm{\theta}}(\bm{z}_{i},t_{i})$ $\triangleright$ learned velocity
+6:  $\bm{z}^{\prime}_{i+1}\leftarrow\bm{z}_{i}+1/T\cdot\bm{v}$ $\triangleright$ discrete integration
+7:  $\bm{z}_{i+1}\leftarrow$ $(\bm{y},\mathcal{A})$-driven update of $\bm{z}^{\prime}_{i+1}$ $\triangleright$ reducing $\ell(\bm{y},\mathcal{A}(\bm{z}))$ starting from $\bm{z}^{\prime}_{i+1}$
 8:end for
@@
 2:Initialize $\bm{z}^{(0)}\sim\mathcal{N}(\bm{0},\bm{I})$
-3:for $e=0$ to $E-1$ do
+3:for $e=0$ to E1E-1 do
 4:  $\bm{z}_{0}\leftarrow\bm{z}^{(e)}$
-5:  for $i=0$ to $T-1$ do $\triangleright$ whole integration path
+5:  for $i=0$ to T1T-1 do $\triangleright$ whole integration path
 6:   $t_{i}\leftarrow i/T$
@@
 9:  end for
-10:  $\bm{z}^{(e+1)}\leftarrow\bm{z}^{(e)}-\eta^{(e)}\cdot\nabla_{\bm{z}}\mathcal{L}(\bm{z}_{i+1})$triangleright reducing loss
+10:  $\bm{z}^{(e+1)}\leftarrow\bm{z}^{(e)}-\eta^{(e)}\cdot\nabla_{\bm{z}}\mathcal{L}(\bm{z}_{i+1})$ $\triangleright$ reducing loss
 11:end for
@@
 2.3 Foundation FM priors for IPs
-Table 1: A comparison between foundation, domain-specific, and untrained priors for Gaussian deblurring on AFHQ (resolution: $\times 256$ 256; DS: domain-specific; FD: foundation).
-Bold: best, & underline: second best, for each metric/column
+Table 1: A comparison between foundation, domain-specific, and untrained priors for Gaussian deblurring on AFHQ-Cat (resolution: $256\times 256$; DS: domain-specific; FD: foundation).
+Bold: best, & underline: second best, for each metric/column.
 […]
-DIP 32.9470 0.9075 0.1873 0.5024
-D-Flow (DS) 30.2512 0.8553 0.1183 0.6536
-D-Flow (FD) 30.0975 0.8861 0.2270 0.5404
-The availability of large-scale training data has enabled the development of foundation generative models for several domains, most of them based on FM (or DM) models, e.g., stable diffusion (v3) [15] and FLUX [14] for images, Sora [22] and Veo for videos, Cosmos world model [16].
-So, recent IP methods based on pretrained FM priors have started to shift from domain-specific priors to these foundation priors.
+DIP 27.5854 0.7179 0.3898 0.2396
+D-Flow (DS) 28.1389 0.7628 0.2783 0.5871
+D-Flow (FD) 25.1453 0.6829 0.5213 0.3228
+FlowDPS (DS) 22.1191 0.5603 0.3850 0.5417
+FlowDPS (FD) 22.1404 0.5930 0.5412 0.2906
+The availability of large-scale training sets has recently fueled intensive development of foundation generative models in several domains, most of them based on FM models and variants, e.g., Stable Diffusion V3 (and newer) [esser2024scalingrectifiedflowtransformers] and FLUX.1 [labs2025flux1kontextflowmatching] for images, OpenAI Sora [OpenAISora2024] and Google Veo [Veo_DeepMind_2025] for videos, and Nvidia Cosmos world model [Agarwal2025Cosmos].
+By contrast, domain-specific FM models are not always readily available (e.g., due to the lack of training data for scientific applications).
+So, recent IP methods based on pretrained FM priors have started to explore foundation priors.
 However, while these foundation priors only constrain the object to be physically meaningful (e.g., the object being a natural image), domain-specific priors provide much more semantic and perhaps structural information about the object (e.g., the object being a facial or brain MRI image), i.e., foundation priors alone are considerably weaker than domain-specific priors.
-In fact, untrained priors such as deep image priors may be powerful enough to promote physically meaningful reconstruction [7, 8, 9, 10, 11].
+In fact, untrained priors such as deep image priors may be powerful enough to promote physically meaningful reconstruction [alkhouri2025understanding, wang2023early, li_deep_2023, zhuang_blind_2023, zhuang_practical_2023].
 A quick comparison summarized in Table 1 confirms our intuition: recent IP methods with foundation priors perform worst compared to domain-specific and even untrained priors.
-To strengthen the object prior, [19] assumes similarity of $\bm{x}$ and $\bm{y}$—valid for typical image restoration problems, and initializes Eq. 2.3 with
+To strengthen the object prior, [ben2024d] assumes similarity of $\bm{x}$ and $\bm{y}$—valid for typical image restoration problems, and initializes Eq. 2.3 with
 $\bm{z}_{0}=\sqrt{\alpha}\bm{y}_{0}+\sqrt{1-\alpha}\bm{z}\quad\text{where}\;\bm{z}\sim\mathcal{N}(\bm{0},\bm{I}),$ (2.4)
-[…]
+where $\bm{y}_{0}$ is the backward solution of the governing ODE, i.e., $\bm{y}_{0}=\bm{y}+\int_{1}^{0}\bm{v}_{\bm{\theta}}(\bm{y}_{t},t)dt$, or inverted seed in other words.
+Moreover, they promote the Gaussianity of the seed $\bm{z}_{0}$ by recognizing that $\norm{\bm{z}_{0}}_{2}^{2}$ follows a $\chi^{2}$ distribution and thus regularizes its negative log-likelihood; [kim2025flowdpsflowdrivenposteriorsampling] takes automatically generated text prompts for $\bm{y}$ from another model as text conditions, as all recent foundation FM models allow text-prompted generation.
 3 Method
-We focus on image restoration with foundation FM priors, and aim to strengthen the image prior based on (1) the similarity of $\bm{x}$ and $\bm{y}$, and (2) Gaussianity in FM priors, in the same vein as [19] but improving upon theirs.
-Adding text prompts as in [17] is presumably helpful, but the reliable generation of high-quality prompts for degraded images seems tricky; we do not pursue it here for leave it for future work.
+We focus on image restoration with foundation FM priors, and aim to strengthen the image prior based on (1) the similarity of $\bm{x}$ and $\bm{y}$, and (2) Gaussianity in FM priors, in the same vein as [ben2024d] but improving upon theirs.
 3.1 Time-dependent warm-up strategy
@@
 One might not worry about this distribution mismatch, as both are supported on the entire ambient space anyway.
-But finite-sample training in practice causes the gap: due to the famous concentration of measure phenomenon [23], virtually all training samples drawn from $\mathcal{N}(\bm{0},\bm{I})$ come from an ultra-thin shell around $\mathbb{S}^{d-1}(\bm{0},\sqrt{d})$, a sphere centered at $\bm{0}$ and with a radius $\sqrt{d}$: $\mathbb{P}[|\norm{\bm{z}}_{2}-\sqrt{d}|\geq t]\leq 2e^{-ct^{2}}$ for a universal constant $c>0$.
+But finite-sample training in practice causes the gap: due to the famous concentration of measure phenomenon [vershynin2018high], virtually all training samples drawn from $\mathcal{N}(\bm{0},\bm{I})$ come from an ultra-thin shell around $\mathbb{S}^{d-1}(\bm{0},\sqrt{d})$, a sphere centered at $\bm{0}$ and with a radius $\sqrt{d}$: $\mathbb{P}[|\norm{\bm{z}}_{2}-\sqrt{d}|\geq t]\leq 2e^{-ct^{2}}$ for a universal constant $c>0$.
 If we write that ultra-thin shell spanned by the finite training set as $\mathcal{S}$, the generation function $\mathcal{G}_{\bm{\theta}}$ is effectively trained on inputs from the domain $\mathcal{S}$, not the entire ambient space: the behavior of $\mathcal{G}_{\bm{\theta}}$ on $\mathcal{S}^{c}$ is largely undetermined.
@@
 Given that the behavior of $\mathcal{G}_{\bm{\theta}}$ can be wild on $\mathcal{S}^{c}$, this initialization strategy seems problematic.
-We take a different route. We know that a typical flow of FM models takes the form:
-$\bm{z}_{t}=\alpha_{t}\bm{x}+\beta_{t}\bm{z}\quad\text{where}\;\bm{z}\sim\mathcal{N}(\bm{0},\bm{I}).$ (3.1)
-Now, when $\bm{x}$ and $\bm{y}$ are close, $\bm{x}=\bm{y}+\bm{\varepsilon}$ for some small $\bm{\varepsilon}$. So, we can write the flow as
-$\bm{z}_{t}=\alpha_{t}(\bm{y}+\bm{\varepsilon})+\beta_{t}\bm{z}\quad\text{where}\;\bm{z}\sim\mathcal{N}(\bm{0},\bm{I})$ (3.2)
-for unknown $\bm{\varepsilon}$.
-When $\alpha_{t}$ is sufficiently small—i.e., we are sufficiently close to $t=0$ in the flow, $\alpha_{t}\bm{\varepsilon}$ can be negligibly small, leading to the approximate flow
-$\bm{z}_{t}\approx\alpha_{t}\bm{y}+\beta_{t}\bm{z}\quad\text{where}\;\bm{z}\sim\mathcal{N}(\bm{0},\bm{I}).$ (3.3)
-In practice, we do not know how small $\alpha_{t}$ should be, so we leave it learnable. This leads to the following formulation
-$)).\displaystyle\boxed{\min\nolimits_{\bm{z},t\in[0,1]}\;\ell\pqty{\bm{y},\mathcal{A}\circ\mathcal{G}_{\bm{\theta}}\pqty{\alpha_{t}\bm{y}+\beta_{t}\bm{z},t}}}.$ (3.4)
+We take a different route. We know that a typical flow of FM models takes the form
+$\bm{z}_{t}=\alpha_{t}\bm{x}+\beta_{t}\bm{z}\quad\text{where}\;\bm{z}\sim\mathcal{N}(\bm{0},\bm{I}),$ (3.1)
+where $\alpha_{t}$ and $\beta_{t}$ are known functions of $t$ with the property that
+$\begin{cases}\alpha_{t}{\scriptstyle\searrow}0,\beta_{t}{\scriptstyle\nearrow}1\;\text{as}\;t\to 0\\ \alpha_{t}{\scriptstyle\nearrow}1,\beta_{t}{\scriptstyle\searrow}0\;\text{as}\;t\to 1,\end{cases}$ (3.2)
+where $\scriptstyle\nearrow$ and $\scriptstyle\searrow$ indicate monotonically increasing and decreasing, respectively.
+Now, when $\bm{x}$ and $\bm{y}$ are close, $\bm{x}=\bm{y}+\bm{\varepsilon}$ for some small (i.e., $\norm{\bm{\varepsilon}}$ is small compared to $\norm{\bm{x}}$ and $\norm{\bm{z}}$) but unknown $\bm{\varepsilon}$.
+So, we can write the flow as
+$\bm{z}_{t}=\alpha_{t}(\bm{y}+\bm{\varepsilon})+\beta_{t}\bm{z}=\alpha_{t}\bm{y}+\beta_{t}\bm{z}+\alpha_{t}\bm{\varepsilon}$
+$\text{where}\;\bm{z}\sim\mathcal{N}(\bm{0},\bm{I}).$ (3.3)
+To eliminate the unknown $\bm{\varepsilon}$, we can approximate the exact flow in Section 3.1 with the following approximate flow
+$\bm{z}_{t}\approx\alpha_{t}\bm{y}+\beta_{t}\bm{z}\quad\text{where}\;\bm{z}\sim\mathcal{N}(\bm{0},\bm{I})$ (3.4)
+with an approximation error $\alpha_{t}\bm{\varepsilon}$.
+To control the error, (1) if $\bm{\varepsilon}$ is relatively large, a small $\alpha_{t}$ is desirable; (2) if $\bm{\varepsilon}$ is already relatively small, a relatively large $\alpha_{t}$ is acceptable.
+So, although we do not know $\bm{\varepsilon}$ itself and hence its magnitude, with appropriate $\alpha_{t}$ we can always make $\alpha_{t}\bm{\varepsilon}$ sufficiently small.
+So, we leave $\alpha_{t}$ as a learnable parameter.
+Since $\alpha_{t}$ is a known function of $t$, we simply need to leave $t\in[0,1]$ learnable, leading to our warm-start formulation
+[…]
 Note that here we overload the notation of $\mathcal{G}_{\bm{\theta}}$ as $\mathcal{G}_{\bm{\theta}}:\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}$—the second input is the current $t$ on the path (the notation we use in Eq.
@@
 Moreover, it is not only grounded in theory and effective in practice (see Section 4), but also speeds up learning as $t>0$ implies shorter flows—although improving the speed is not our current focus.
-Due to the approximation errors in matching the ideal flow during FM training, as well as when approximating Eq.
-3.2 using Eq.
-3.3, the distribution of $\bm{z}_{t}$ could be slightly off the ideal distribution.
+Due to the approximation errors in matching the ideal flow during FM training, as well as when approximating Section 3.1 using Eq.
+3.4, the distribution of $\bm{z}_{t}$ could be slightly off the ideal distribution.
 To rectify this, we perform a lightweight variance calibration in our implementation:
 $\hat{\bm{z}}_{t}=\sqrt{\mathrm{Var}(Z_{t})/\mathrm{Var}(\bm{z}_{t})}\cdot\bm{z}_{t},$ (3.6)
-where $\mathrm{Var}(Z_{t})$ is the scalar-variable variance across all dimensions and the generation paths of 512512512 samples, and $\mathrm{Var}(\bm{z}_{t})$ scalar-variable variance for $\bm{z}_{t}$ across all dimensions.
+where $\mathrm{Var}(Z_{t})$ is the scalar-variable variance across all dimensions and the generation paths of $512$ samples, and $\mathrm{Var}(\bm{z}_{t})$ scalar-variable variance for $\bm{z}_{t}$ across all dimensions.
 3.2 Gaussianity regularization
@@
 If $\bm{z}_{0}\sim\mathcal{N}(\bm{0},\bm{I})$, $\norm{\bm{z}_{0}}_{2}^{2}\sim\chi^{2}(d)$ and the negative log-likelihood is $h(\bm{z}_{0})=-(d/2-1)\log\norm{\bm{z}_{0}}_{2}^{2}+\norm{\bm{z}_{0}}_{2}^{2}/2+C$ for some constant $C$ independent of $\bm{z}_{0}$.
-[19] promotes the Gaussianity of $\bm{z}_{0}$ by regularizing $h(\bm{z}_{0})$.
+[ben2024d] promotes the Gaussianity of $\bm{z}_{0}$ by regularizing $h(\bm{z}_{0})$.
 While $h(\bm{z}_{0})$ is minimized at any $\bm{z}_{0}$ satisfies $\norm{\bm{z}_{0}}_{2}^{2}=d-2$, away from this value the function changes slowly; see Fig.
@@
 For our formulation in Eq.
-3.4, we hope to promote the Gaussianity of $\bm{z}$.
+3.5, we hope to promote the Gaussianity of $\bm{z}$.
 To enforce the sharp concentration of $\bm{z}$, we project $\bm{z}$ onto the sphere $\mathbb{S}^{d-1}(\bm{0},\sqrt{d})$ in each step by
 $\bm{z}^{\prime}=\sqrt{d}\bm{z}/\|\bm{z}\|_{2}.$ (3.7)
-We are not the first to use the more precise spherical constraint or regularization in promoting Gaussianity in the DM/FM literature; see, e.g., [24].
+We are not the first to use the more precise spherical constraint or regularization in promoting Gaussianity in the DM/FM literature; see, e.g., [yang2024guidancesphericalgaussianconstraint].
 4 Experiment
 Table 2: x4 Super Resolution with additive Gaussian noise ($\sigma=0.03$). (Bold: best, underline: second best)
-DIV2K [25] AFHQ [26]
+DIV2K [Agustsson_2017_CVPR_Workshops] AFHQ [choi2020starganv2]
 […]
@@
 Table 3: Gaussian Deblur with additive Gaussian noise ($\sigma=0.03$). (Bold: best, underline: second best)
-DIV2K [25] AFHQ [26]
+DIV2K [Agustsson_2017_CVPR_Workshops] AFHQ [choi2020starganv2]
 […]
@@
 We test our FMPlug method against other SOTA methods on two linear IPs: super-resolution and Gaussian debluring.
-Datasets, tasks, and evaluation metrics  Following [17], we use two datasets: 252525 DIV2K  [25] validation images and 252525 AFHQ cat  [26] images.
-We set the image resolution to $\times 512$ 512 by resizing and cropping the original.
-We consider two tasks: i) 4$\times$ super-resolution (SR) from $\times 128$ 128 to $\times 512$ 512; ii) Gaussian deblurring (GD) with a kernel size of 616161 and standard deviation of 3.03.03.0.
-Both tasks are added with Gaussian noise $\sigma=0.03$.
-For metrics, we use PSNR and SSIM for pixel-level difference, and LPIPS for perceptual difference, and CLIPIQA, a no-reference quality metric.
 […]
+Competing methods  We compare our FMPlug (W: warm-up, R: regularized, Number of Function Evaluations (NFE) =3=3) with Deep Image Prior [Ulyanov_2020] (an untrained image prior), D-Flow ($\text{NFE}=3$) [ben2024d] (a SOTA plug-in method), FlowDPS ($\text{NFE}=28$) [kim2025flowdpsflowdrivenposteriorsampling] (a SOTA interleaving method) and FlowChef ($\text{NFE}=28$) [patel2024steering] (another SOTA interleaving method).
+For a fair comparison, we use the pretrained SD3 model [esser2024scalingrectifiedflowtransformers] as the backbone for all methods that require pretrained priors, i.e., foundation priors.
 No text-prompt For a fair comparison, we disable text prompt and classifier-free guidance for all methods. It is also not clear how reliably we can generate pertinent text prompts for degraded images.
@@
 This confirms the benefits brought about by the precise Gaussianity regularization in Eq.
-3.6.
+3.7.
 (a) Measurement
@@
 (h) FMPlug-W-R
-Fig. 2: Qualitative results for 4×4$\times$4 × super-resolution on the AFHQ dataset.
+Fig. 2: Qualitative results for $4\times$ super-resolution on the AFHQ dataset.
 (a) Measurement
@@
 (h) FMPlug-W-R
-Fig. 3: Qualitative results for 4×4$\times$4 × super-resolution on the DIV2K dataset.
+Fig. 3: Qualitative results for $4\times$ super-resolution on the DIV2K dataset.
 (a) Measurement